What concentration of SO2−3 is in equilibrium with Ag2SO3(s) and 4.60×10−3 M Ag+ ?
1
Expert's answer
2020-04-22T09:17:43-0400

Ag2SO3= 2Ag+ + SO3 2-

C(SO3 2-)= C(Ag+)/2=2.3*10-3